OpenStudy (anonymous):

Given the function g(x) = −2x + 4, find g(−5).

OpenStudy (anonymous):

g[-5]=14

OpenStudy (anonymous):

replace x by -5 get \[g(-5)=-2\times -5+4=10+4=14\]